An electric lamp of 100 Ω, a toaster of resistance. 50 Ω , and a water filter of resistance 500 Ω are connected in parallel to a 220 V source. What is the resistance of an electric iron connected to the same source that takes as much current as all the three appliances, and what is the current through it?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Resistance of electric lamp =100 ohm
Resistance of a toaster = 50 ohm
Resistance of a water filter = 500 ohm
Potential difference = 220 volt
Total resistance of the three appliances in parallel = 1/R = 1/R1 + 1/R2 + 1/R3
= 1/100 + 1/50 + 1/500
1/R = 5 + 10 + 1/500
1/R = 16/500 = 4/125
= 125/4 = 31.25 ohm
V = I × R
220 = I × 31.25
I = 7.04 amp
Hence current through the electric iron = 7.04 amp.

Thanks